OpenStudy (onepieceftw):

Can someone help me solve this problem involving trig identities? Problem in comments.

OpenStudy (onepieceftw):

\[\frac{ \cos^2 \theta }{ \sin^2 \theta } + \csc \theta \sin \theta\]

OpenStudy (solomonzelman):

Let's write what each of these means when finding sides of a right triangle. \(\color{#000000}{ \displaystyle \sin \theta =({\rm opposite / hypotenuse}) }\) \(\color{#000000}{ \displaystyle \cos \theta =({\rm adjacent / hypotenuse}) }\) And you know that: \(\color{#000000}{ \displaystyle \tan \theta = \frac{\rm opposite}{\rm adjacent} }\) Therefore, when you divide top and bottom by adjacent you get: \(\color{#000000}{ \displaystyle \tan \theta = \frac{({\rm opposite / hypotenuse})}{({\rm adjacent / hypotenuse})} }\) And the substitution yields, \(\color{#000000}{ \displaystyle \tan \theta = \frac{ \sin \theta}{\cos \theta} }\)

OpenStudy (solomonzelman):

And the \(\color{#000000}{ \displaystyle \csc \theta = \frac{ 1}{\sin \theta} }\), because the \(\color{#000000}{ \displaystyle \csc\theta }\) is the reciprocal of the \(\color{#000000}{ \displaystyle \sin \theta }\).

OpenStudy (solomonzelman):

You can further imply that: \(\color{#000000}{ \displaystyle \tan \theta = \frac{ \sin \theta}{\cos \theta} }\) \(\color{#000000}{ \displaystyle \left(\tan \theta \right)^2= \left(\frac{ \sin \theta}{\cos \theta} \right)^2 }\) \(\color{#000000}{ \displaystyle \tan^2\theta= \frac{ \sin^2\theta}{\cos^2 \theta} }\)

OpenStudy (solomonzelman):

And, \(\color{#000000}{ \displaystyle \sin\theta\times \csc\theta }\) is just a product of a number and its reciprocal, which you should equal 1.

OpenStudy (solomonzelman):

Then, you will need another rule: \(\color{#000000}{ \displaystyle \tan^2\theta+1=\sec^2\theta }\) ----------------------------

OpenStudy (onepieceftw):

I was thinking that because it is cos/sin it would be equal to cot because cot = cos/sin Am I wrong?

OpenStudy (onepieceftw):

So then \[\cot^2 + 1 = \csc^2\] which is what wolfram alpha says it is.

OpenStudy (onepieceftw):

Regardless if you or I had a faux pas I figured it out with your help. Thank you very much!

OpenStudy (solomonzelman):

No, you are totally correct.

OpenStudy (solomonzelman):

Nice site to refer to :) In fact, I am using Mathematica program for some of tutorials/projects, and that is also really useful for everything in math like matrix reduction or almost anything yo can think of.

OpenStudy (solomonzelman):

yw